# Dejavu: Towards Experience Feedback Learning for Embodied Intelligence

### v2

@@ E.2.2 Integration Cost
E.3 Residual correction vs. direct action prediction
@@
$\mathbf{a}_{t}\;=\;\mathbf{a}^{\text{base}}_{t}\;+\;\Delta\mathbf{a}^{\text{EFN}}_{t},$ (48)
@@
$\mathbf{a}_{t}\;=\;\mathbf{a}^{\text{EFN}}_{t},$ (49)
using the same frozen VLA features and the same lightweight MLP head as in our residual formulation (see Tab. 4 for the architectural footprint). All other training hyperparameters (SAC configuration, reward shaping, and interaction budget) are kept identical to the residual version.
Table 9: Residual vs. direct-action EFN on LIBERO Environment. The direct-action variant fails to learn and remains close to zero.
@@
Empirically, this direct-action EFN fails to learn meaningful behavior in our environments: across both simulation environments and real-world platform, the success rate remains close to zero throughout training (as shown in Table 9) and the learning curves exhibit strong oscillations instead of stable improvement. We thus omit this variant from the main tables and only report it here as a negative result. We hypothesize two main reasons. First, by removing the base policy we effectively discard the generalization ability acquired by the pre-trained VLA, forcing a small EFN head to re-learn the entire control policy from sparse rewards and high-dimensional observations. Second, the head is deliberately designed to be extremely lightweight to meet our latency budget; while this is sufficient for modeling residual corrections around a strong backbone, it appears under-parameterized for modeling the full continuous action manifold from scratch. Taken together, these findings justify our choice of treating EFN as a residual corrector rather than a standalone policy.
@@ Appendix F Visualization
To complement the quantitative results in Table 2, we provide real-world demonstration videos recorded on the AgiBot-G1 manipulation platform in the demo/ directory. For each real-world task evaluated in Table 2, we include a video of the EFN-augmented policy executing the task (e.g., StockLift_with_EFN.mp4), illustrating how the deployed agent uses retrieved experiences to refine the base VLA policy and produce more reliable executions.
Due to space limitations in the supplementary material, we provide one paired comparison (with vs. without EFN) for the most challenging task, DrawerStore, whose baseline success rate without EFN is only 5.3% while EFN raises it to 58.7% (see Table 2). The video DrawerStore_without_EFN.mp4 was obtained after nearly twenty recording attempts, reflecting the low baseline success probability; even in this successful rollout, the robot exhibits hesitant behavior, including repeated grasps and an imprecise final release that almost leads to failure. In contrast, DrawerStore_with_EFN.mp4 was recorded easily and shows a smooth, near-perfect execution of the same task, with a decisive grasp, accurate drawer interaction, and stable object placement. This qualitative gap is consistent with the quantitative improvement brought by EFN and provides an intuitive illustration of how experience feedback stabilizes behavior on the hardest real-world task.

### v3

@@ E.1.2 Integration Cost
E.2 Residual correction vs. direct action prediction
@@
$\mathbf{a}_{t}\;=\;\mathbf{a}^{\text{base}}_{t}\;+\;\Delta\mathbf{a}^{\text{EFN}}_{t},$ (E.1)
@@
$\mathbf{a}_{t}\;=\;\mathbf{a}^{\text{EFN}}_{t},$ (E.2)
using the same frozen VLA features and the same lightweight MLP head as in our residual formulation (see Tab. B.1 for the architectural footprint). All other training hyperparameters (SAC configuration, reward shaping, and interaction budget) are kept identical to the residual version.
Table E.3: Residual vs. direct-action EFN on the LIBERO benchmark. The direct-action variant fails to learn and remains close to zero.
@@
Empirically, this direct-action EFN fails to learn meaningful behavior in our environments: across both simulation and the real-world platform, the success rate remains close to zero throughout training (as shown in Table E.3), and the learning curves exhibit strong oscillations rather than stable improvement. We therefore omit this variant from the main tables and report it here only as a negative result. We hypothesize two main reasons. First, by removing the base policy we effectively discard the generalization ability acquired by the pre-trained VLA, forcing a small EFN head to re-learn the entire control policy from sparse rewards and high-dimensional observations. Second, the head is deliberately designed to be extremely lightweight to meet our latency budget; while this is sufficient for modeling residual corrections around a strong backbone, it appears under-parameterized for modeling the full continuous action manifold from scratch. Taken together, these findings justify our choice of treating EFN as a residual corrector rather than a standalone policy.
@@ Appendix F Visualization
To complement the quantitative results in Table 2 in the main paper, we provide real-world demonstration videos recorded on the AgiBot-G1 manipulation platform in the demo/ directory. For each real-world task evaluated in Table 2 in the main paper, we include a video of the EFN-augmented policy executing the task (e.g., StockLift_with_EFN.mp4), illustrating how the deployed agent uses retrieved experiences to refine the base VLA policy and produce more reliable executions.
Due to space limitations in the supplementary material, we provide one paired comparison (with vs. without EFN) for the most challenging task, DrawerStore, whose baseline success rate without EFN is only 5.3% while EFN raises it to 58.7% (see Table 2 in the main paper). The video DrawerStore_without_EFN.mp4 was obtained after nearly twenty recording attempts, reflecting the low baseline success probability; even in this successful rollout, the robot exhibits hesitant behavior, including repeated grasps and an imprecise final release that almost leads to failure. In contrast, DrawerStore_with_EFN.mp4 was recorded easily and shows a smooth, near-perfect execution of the same task, with a decisive grasp, accurate drawer interaction, and stable object placement. This qualitative gap is consistent with the quantitative improvement brought by EFN and provides an intuitive illustration of how experience feedback stabilizes behavior on the hardest real-world task.